OpenStudy (anonymous):

2x+6y=0 Find the constant of variation k for the direct variation.

Vocaloid (vocaloid):

start by subtracting 2x from each side

OpenStudy (anonymous):

do i want to find x or y

Vocaloid (vocaloid):

neither, just do what I asked

OpenStudy (anonymous):

i did

Vocaloid (vocaloid):

like I said, subtract 2x from each side and tell me what you get

Vocaloid (vocaloid):

2x+6y-2x = ?

OpenStudy (anonymous):

when you subtract 2x from both sides you get 6y=-2x

Vocaloid (vocaloid):

good! now divide both sides by 6

OpenStudy (anonymous):

so -1/3?

Vocaloid (vocaloid):

yeah, k = -1/3 since y = (-1/3)x

OpenStudy (anonymous):

okay thank you